# CHESS: Contextual Harnessing for Efficient SQL Synthesis

### v1

@@ Abstract
Utilizing large language models (LLMs) for transforming natural language questions into SQL queries (text-to-SQL) is a promising yet challenging approach, particularly when applied to real-world databases with complex and extensive schemas. In particular, effectively incorporating data catalogs and database values for SQL generation remains an obstacle, leading to suboptimal solutions. We address this problem by proposing a new pipeline that effectively retrieves relevant data and context, selects an efficient schema, and synthesizes correct and efficient SQL queries. To increase retrieval precision, our pipeline introduces a hierarchical retrieval method leveraging model-generated keywords, locality-sensitive hashing indexing, and vector databases. Additionally, we have developed an adaptive schema pruning technique that adjusts based on the complexity of the problem and the model’s context size. Our approach generalizes to both frontier proprietary models like GPT-4 and open-source models such as Llama-3-70B. Through a series of ablation studies, we demonstrate the effectiveness of each component of our pipeline and its impact on the end-to-end performance. Our method achieves new state-of-the-art performance on the cross-domain challenging BIRD dataset.
@@ 1 Introduction
Translating natural language questions into database queries, or text-to-SQL, is a long-standing research problem. This issue has been exacerbated in recent years due to the growing complexity of databases, driven by the increasing sizes of schemas (sets of columns and tables), values (content), and catalogs (metadata describing schemas and values) stored within them. Even the largest proprietary models, such as GPT-4, lag significantly behind human performance on text-to-SQL benchmarks, with a notable accuracy gap of 30% (Li et al., 2024b). Beyond the complexity of writing SQL queries, this substantial gap is primarily caused by the need to effectively retrieve and integrate multiple sources of information, including database values, catalogs, and schema, each in different formats, which complicates the process.
@@
In Figure 1, we show some of the challenges facing modern text-to-SQL systems. For instance, users’ questions might not directly match the stored values in the database, making it crucial to accurately identify the value format for effective SQL query formulation. Additionally, real-world database schemas often contain ambiguous column names, table names, and messy data, complicating the SQL translation process and necessitating a robust retrieval system to identify relevant information. Moreover, there are typically multiple valid SQL queries that could answer the same question. For example, for the question illustrated on the right side of the Figure 1, one might use ’ORDER BY’ and ’LIMIT 1’ to find the highest average score, while another approach could involve a subquery with the ’MAX()’ function, potentially leading to different outputs.
Figure 1: Example of challenges in text-to-SQL translation. 1) Questions passed by the users might not have the exact database value. 2) Column names might not be a good representation of what they store so using database catalogs is an essential part of text-to-SQL translation. 3) For a given question there are multiple ways of writing a correct SQL answer.
Earlier work in the area (Pourreza and Rafiei, 2024a; Wang et al., 2023; Qi et al., 2022; Rajkumar et al., 2022; Li et al., 2024c) has generally limited the context for SQL generation to table structures, column definitions, and sample rows. However, in production-level databases, the database catalog and database values constitute a rich source of information that is crucial for generating accurate SQL queries.
We introduce “CHESS: contextual harnessing for efficient SQL synthesis", an end-to-end text-to-SQL system that targets real-world and complex databases. CHESS introduces a scalable and effective LLM-based pipeline for SQL generation that consists of three main components: entity and context retrieval, schema selection, and SQL generation.
@@
For entity and context retrieval, we present scalable and efficient methods using locality-sensitive hashing to retrieve database values from millions of rows, leveraging keyword detection, and vector databases to extract contextual information from database catalogs. Our approach utilizes both semantic and syntactic similarities between the database content and the user’s query to enhance SQL prediction accuracy. In the schema selection phase, we utilize the retrieved information to narrow down the initial schema with potentially hundreds of columns to an efficient set of columns, usually less than ten. Throughout this step, we extract a minimal yet sufficient subset of the database schema. Finally, the extracted database schema is passed to a query generation module, which uses our fine-tuned SQL generator model combined with a revision step to effectively generate a SQL query.
Through a series of ablation studies 4.4, we demonstrate the critical role of each module in the pipeline in guiding LLMs to generate accurate SQL queries. Specifically, our entity and context retrieval module contributes substantially to performance, as evidenced by a 5% accuracy improvement.
At the time of submission, CHESS ranks first among all disclosed methodologies on BIRD (Li et al., 2024b), with a 65% and 66.69% execution accuracy on the development and test set respectively. To our knowledge, BIRD is the most challenging real-world text-to-SQL benchmark that is publicly available. It features more than 12000 unique question-SQL pairs, spanning 37 professional fields, including healthcare, education, blockchain, and sports, and covering 95 large databases with a combined size of 33.4 GB. An active leaderboard for BIRD is maintained by a third party (with a private test set), which we used to evaluate the performance of CHESS. CHESS also ranks second among all methods, with a marginal gap to the propriety (and undisclosed) approach that currently has a test set accuracy of 67.86% on BIRD’s leaderboard.
We also provide an end-to-end open-source version of CHESS that obtains the best performance among other open-source baselines on the BIRD development set, with an execution accuracy of 61.5%. The use of open-source models for text-to-SQL is crucial, particularly when databases may contain private information that should not be shared with third-party LLM providers.
Concretely, our contributions are as follows:
  • Breaking down the Text-to-SQL task into a 3-staged pipeline, including entity and context retrieval, schema selection, and query generation
  • A scalable hierarchical retrieval approach for extracting the important entities and contexts
  • An efficient three-staged schema pruning protocol, consisting of individual column filtering, table selection, and a final column selection for extracting a minimally sufficient schema
  • A fine-tuned open-source DeepSeek-33B Coder model with a novel training dataset construction approach with noise injection to mitigate error propagation
  • A high-performing end-to-end open-source pipeline ensuring the privacy of the information
  • Setting new state-of-the-art results on the BIRD dataset among known methodologies
@@ 2 Related Work
Generating accurate SQL queries from natural language questions, known as text-to-SQL is an active area of research within both the natural language processing (NLP) and database communities. Early efforts by the database community approach the problem through custom templates (Zelle and Mooney, 1996) marked initial advancements in this field, albeit at the expense of significant manual effort. Recently, text-to-SQL methodologies have increasingly incorporated transformer-based models, particularly sequence-to-sequence architectures (Vaswani et al., 2017; Sutskever et al., 2014). These sequence-to-sequence models, capable of end-to-end training, are particularly well-suited for tasks requiring the generation of one sequence from another, such as translation, summarization, and text-to-SQL (Qin et al., 2022).
[…]
@@
Alongside the widespread adoption of LLMs across various NLP domains, the text-to-SQL field has similarly benefited from recent methodological innovations with LLMs to enhance performance. Early approaches (Rajkumar et al., 2022), leveraged the zero-shot in-context learning capabilities of LLMs for SQL generation. Building on this, subsequent models including DIN-SQL (Pourreza and Rafiei, 2024a), DAIL-SQL (Gao et al., 2023), MAC-SQL (Wang et al., 2023), and C3 (Dong et al., 2023) have enhanced LLM performance through task decomposition and techniques such as Chains of Thought (CoT) (Wei et al., 2022), self-consistency (Wang et al., 2022), and least-to-most prompting (Zhou et al., 2022). In addition to in-context learning, proposals in DAIL-SQL (Gao et al., 2023), DTS-SQL (Pourreza and Rafiei, 2024b), and CodeS (Li et al., 2024a) have sought to elevate the capabilities of open-source LLMs through supervised fine-tuning, aiming to rival or exceed their larger, proprietary counterparts. However, the most notable performance gains have been observed in proprietary LLMs utilizing in-context learning methods (Li et al., 2024b). Unlike previous efforts, this paper introduces a hybrid approach that combines both in-context learning and supervised fine-tuning to further enhance performance. Moreover, we propose novel methods to integrate contextual data such as database values and database catalog into the text-to-SQL pipeline, leveraging a rich yet often overlooked source of information.
Independently, but concurrent with our work, MCS-SQL (Lee et al., 2024) introduced a method that relies on using multiple prompts and sampling several responses from LLMs to mitigate sensitivity to the arrangement of tables, columns, and few-shot in-context samples. They also devised a technique to filter out irrelevant table and column names. However, unlike our approach, they do not emphasize retrieving pertinent information from the database catalog and database values. Instead, they rely on using a large number of samples from LLMs to reduce prompt sensitivity for accuracy improvement.
@@ 3 Methodology
The text-to-SQL task is a challenging learning problem, requiring the model to translate a natural language question and database schema into a valid SQL query. The input consists of the text of the question, the database schema defining table structures and column types, and the database instance containing the actual or samples of the table content. Further, the input may include a database catalog with metadata, allowing free-form textual attribute values. Key challenges include noisy database content, the need to implicitly capture semantic correspondences between input elements, and the compositional nature of mapping language to SQL’s formal query structure.
"CHESS: Contextual Harnessing for Efficient SQL Synthesis" is an end-to-end text-to-SQL system designed for complex, real-world databases. It introduces a scalable and effective LLM-based pipeline consisting of three main components: entity and context retrieval, schema selection, and SQL generation. The entity and context retrieval component 3.1 uses keyword selection, locality-sensitive hashing, and vector databases to efficiently retrieve relevant database values and contextual information from large databases. The schema selection phase 3.2 then narrows down the initial schema to a minimal yet sufficient subset of columns. Finally, the extracted schema is passed to a query generation module 3.3 that utilizes a fine-tuned SQL generator model and a revision step to generate an accurate SQL query. The overall pipeline is demonstrated in Figure 2.
Finally, some of the implementation details of our method, including how we preprocess database values and data catalogs to expedite retrieval during the pipeline execution, are provided in Appendix C. Detailed execution traces of our pipeline are presented in Appendix E.
Figure 2: Our pipeline with modules for entity and context retrieval, schema selection, and query generation.
A key feature of our SQL query generation approach is the integration of multiple sources of information, including database catalogs, values, and schemas. However, these datasets with schemas containing hundreds of columns, each with detailed descriptions, and potentially millions of rows are often very large. Passing all this information to an LLM is often impractical due to limited context windows. Even if feasible, it can negatively impact the LLM’s reasoning capabilities, as demonstrated in (Hsieh et al., 2024) and our own ablation studies in Section 4.5. Our pipeline addresses this challenge by providing the LLM with minimal yet sufficient information necessary for each task. Maintaining minimal sufficiency is a key feature in all modules of the pipeline. Most crucially, during the SQL generation phase, we try to identify and pass to the model only the columns that are needed for the generation of the SQL query.
3.1 Entity and Context Retrieval
The first module in the pipeline identifies the relevant information in the input, including the entities referred to in the question and the contextual information provided about them in the database schema. This is done in three steps.
Keyword Extraction.
To search for similar values in the database and schema descriptions, we first need to extract the main keywords from the natural-language question. Our approach to this problem is to prompt the model with few-shot examples of the task and the question, asking it to identify and extract keywords, keyphrases, and named entities.
Entity Retrieval.
@@ 3.1 Entity and Context Retrieval
From the list of keywords extracted from the question, some may correspond to entities present in the database values. In this step, we search for similar values in the database and return the most relevant ones, along with their corresponding columns, for each keyword. As illustrated in Figure 1, searching for exact matches of the keywords cannot handle variations or typos, necessitating a more flexible search method. To measure the syntactic similarity between the keywords and the database values, we use the edit distance similarity metric. Additionally, to make the retrieval process more efficient, we propose a hierarchical retrieval strategy based on Locality Sensitive Hashing (LSH) and semantic (embedding) similarity measures, which we explain in detail in Appendix C. This approach allows us to efficiently retrieve values that exhibit a high degree of both syntactic and semantic similarity to the keywords.
Context Retrieval.
@@
In addition to the values, database catalogs explaining the schema may be available. For instance, each column may have a description, an extended column name (in the case of abbreviations), and a value description. As shown in Figure 1, this information can be useful, and not providing it to the model can lead to suboptimal performance. As explained before, in retrieving this context, we aim to identify only the minimally sufficient or the most relevant information. This is done by retrieving the most similar descriptions to the extracted keywords, measured by a semantic (embedding) similarity metric when querying the vector database of descriptions created during the preprocessing step.
3.2 Schema Selection
Our goal in this step is to narrow down the schema to include only the necessary tables and columns needed for generating the SQL query. We refer to this optimized set of necessary tables and columns as the efficient schema. Achieving an efficient schema leads to better performance in SQL query generation by excluding irrelevant information. We use recall and precision metrics to determine whether we have selected the correct tables and columns using the correct SQL query as the ground truth. The results are presented in Table 5 along with a detailed example of the schema selection process in Appendix F. The prompts for all of the sub-modules in schema selection are augmented with chain-of-though prompting Wei et al. (2022) to improve the reasoning ability of LLMs which is essential for this task.
Individual Column Filtering.
A database can contain hundreds of columns, many of which may be semantically irrelevant to the question. Looking for an efficient schema, we aim to filter out the irrelevant columns and pass only the most relevant ones to the table selection step. To accomplish this, we treat the relevance of each column to the question as a binary classification task for the model, essentially asking the LLM if the column may be relevant to the question. This step is useful only for removing the irrelevant columns that are obvious, but evaluating the relevance of a column in isolation is not always possible. We address this limitation in the subsequent steps, table selection and column selection, in which we give the model a more global view of the schema.
Table Selection.
After filtering out irrelevant columns, we proceed to select the tables essential for generating the SQL query. In this step, we present the model with the filtered schema from the previous step and ask it to assess the relevance of each table, selecting only those necessary for the SQL query.
Final Column Selection.
In the final step of schema selection, we aim to reduce the schema to the minimal set of columns necessary for generating the SQL query. We prompt the model to evaluate the necessity of each column in the filtered tables. This step includes a chain-of-thought explanation of why each column is needed, followed by the selection of the required columns.
3.3 Query Generation
At this point, we have selected an efficient schema augmented by the relevant context, containing all the necessary information to craft a SQL query that answers the question. In the following steps, we first write a candidate SQL query and then revise it to fix potential semantic and syntactic errors.
Candidate Generation.
After reducing the schema to the minimal set of tables and columns, we prompt the model to generate an SQL query that answers the question. In the prompt, we provide the minimal schema obtained from the previous steps, along with the relevant values and descriptions retrieved in the first step of the pipeline. With this information, the model generates a candidate SQL query.
Revision.
In the final step of the pipeline, we aim to fix potential logical and syntactic errors in the candidate SQL query. We provide the model with the database schema, the question, the generated candidate SQL query, and its execution result. The model is then asked to evaluate the correctness of the SQL query and revise it if necessary.
To assist the model in identifying and correcting mistakes, we give it a set of rules following (Bai et al., 2022). In some cases, there may be multiple ways to correct the candidate SQL query. Even with zero-temperature sampling, the model might output different corrections across multiple samplings. To reduce the noise in the model’s output, we use self-consistency to select the SQL query that appears most consistently across three samples.

### v3

@@ Abstract
Translating natural language questions into SQL queries, known as text-to-SQL, is a long-standing research problem. Effective text-to-SQL synthesis can become very challenging due to (i) the extensive size of database catalogs (descriptions of tables and their columns) and database values, (ii) reasoning over large database schemas, (iii) ensuring the functional validity of the generated queries, and (iv) navigating the ambiguities of natural language questions. We introduce CHESS, a Large Language Model (LLM) based multi-agent framework for efficient and scalable SQL synthesis, comprising four specialized agents, each targeting one of the aforementioned challenges: the Information Retriever (IR) extracts relevant data, the Schema Selector (SS) prunes large schemas, the Candidate Generator (CG) generates high-quality candidates and refines queries iteratively, and the Unit Tester (UT) validates queries through LLM-based natural language unit tests. Our framework offers configurable features that adapt to various deployment constraints, including 1) Supporting industrial-scale databases: leveraging the Schema Selector agent, CHESS efficiently narrows down very large database schemas into manageable sub-schemas, boosting system accuracy by approximately 2% and reducing the number of LLM tokens by ×$\times 5$. 2) State-of-the-Art privacy-preserving performance: Among the methods using open-source models, CHESS achieves state-of-the-art performance, resulting in a high-performing, privacy-preserving system suitable for industrial deployment. 3) Scalablity with additional compute budget: In settings with high computational budgets, CHESS achieves 71.10% accuracy on the BIRD test set, within 2% of the leading proprietary method, while requiring approximately 83% fewer LLM calls.
footnotetext: 1Stanford University, Stanford, CA, USA 2University of Alberta, Alberta, Canada. Correspondence to: Shayan Talaei <stalaei@stanford.edu>.
11footnotetext: Equal senior authorship.
@@ 1 Introduction
Generating SQL queries from natural language questions, commonly known as text-to-SQL, is a persistent and high-impact research problem. The growing complexity of databases in recent years has exacerbated this problem, due to the increasing sizes of schemas (sets of columns and tables), values (content), and catalogs (metadata describing schemas and values) stored within them. Even some of the largest proprietary models, such as GPT-4, lag significantly behind human performance on text-to-SQL benchmarks, with a notable accuracy gap of 40% (Li et al., 2024b). Beyond the complexity of writing SQL queries, this substantial gap is primarily caused by the need to effectively retrieve and integrate multiple sources of information, including database values, catalogs, and schema, each in different formats, which complicates the process. Considering all these complexities, the lack of a robust verification method during the inference stage of text-to-SQL systems can undermine the reliability of the generated SQL queries.
@@
Figure 1 illustrates some of the challenges facing modern text-to-SQL systems. For instance, users’ questions might not directly match the stored values in the database (Gan et al., 2021), making it crucial to accurately identify the value format for effective SQL query formulation. Additionally, real-world database schemas often contain ambiguous column names, table names, and messy data, further complicating the SQL translation process and necessitating a robust retrieval system to identify relevant information (Pourreza and Rafiei, 2023). Moreover, some questions may be inherently ambiguous, leading large language models (LLMs) to make subtle mistakes when generating SQL queries. For example, in the question illustrated on the right side of Figure 1, one approach might use ORDER BY and LIMIT 1 to identify the highest average score, thereby returning only a single result. Another approach could involve a subquery with the MAX() function, which would return all entries that match the maximum AvgScrRead value. This distinction in methods can lead to different outputs.
Figure 1: Examples of challenges in text-to-SQL translation. (1) Users’ questions might not have the exact database value. (2) Column names might not accurately represent their content, making database catalogs essential. (3) Ambiguity in question interpretation can result in different, seemingly valid SQL queries.
CHESS Framework. To address these challenges, we introduce CHESS: Contextual Harnessing for Efficient SQL Synthesis, a multi-agent framework designed to improve SQL query generation in complex, real-world databases. Each agent within CHESS is specifically designed to tackle one or more of the aforementioned challenges:
@@
The Information Retriever (IR) agent aims to retrieve the relevant entities (values in the database) and context (schema descriptions provided in the database catalog). To achieve this, we present scalable and efficient methods using locality-sensitive hashing to retrieve database values from millions of rows, leveraging keyword detection, and vector databases to extract contextual information from database catalogs. Our approach utilizes both semantic and syntactic similarities between the database content and the user’s query to enhance the retrieval quality.
As previous works (Guerreiro et al., 2022) and our ablation studies (Figure 3) confirm, overloading the LLM context with redundant information can degrade its performance on downstream task. To address this, the Schema Selector (SS) agent is specifically designed to prune very large database schemas, allowing the candidate generator agent (CG) to perform a more effective reasoning when synthesizing SQL queries. This SS agent is equipped with three tools, filter_column, select_tables, and select_columns, which allow it to maintain a scalable performance even with complex, real-world schemas with more than 4000 columns.
The Candidate Generator (CG) agent generates SQL queries based on the question posed to the database. It then executes and revises these queries as needed based on the results.
The Unit Tester (UT) agent assesses the quality of the final query by generating multiple natural language unit tests and evaluating the candidate queries against them. UT assigns a score to each candidate based on passed tests, selecting the highest-scoring query, thus ensuring robust query generation.
In summary, the CHESS framework leverages a collaborative multi-agent approach to enhance SQL generation for complex databases. Each agent contributes a specialized function, from retrieving relevant data and pruning schema information to generating and validating SQL queries, collectively ensuring accurate, efficient, and context-aware query synthesis.
CHESS’s Adaptivity for Deployment Constraints. Real-world deployment of text-to-SQL systems must consider various constraints, such as the reasoning capabilities and sampling budgets of large language models (LLMs), as well as data privacy concerns. CHESS is designed as a configurable framework that can adapt to different deployment scenarios while maintaining high performance. In subsection 4.2, we discuss various deployment settings and how CHESS can be customized to meet performance requirements while adhering to these constraints.
To address the challenge of complex industrial database schemas, CHESS incorporates a Schema Selector (SS) agent that efficiently narrows down large schemas into manageable sub-schemas, ensuring sufficient information to accurately answer queries. For environments with limited computational resources, CHESS optimizes performance by minimizing reliance on frequent and costly LLM calls, achieving comparable results using less resource-intensive models. Regarding data privacy concerns, CHESS supports the use of fully open-source models that can be deployed on-premise, eliminating the need to send private data to third-party providers. These features collectively enable CHESS to adapt to various deployment constraints, offering a flexible and high-performing solution for industrial text-to-SQL applications.
Experiments. To validate the effectiveness of CHESS, we conducted comprehensive experiments on challenging benchmarks, including BIRD and Spider, across various deployment scenarios. Our evaluations included testing on large-scale schemas, operating under limited computational resources, and ensuring data privacy by using fully open-source models. We also performed ablation studies to assess the contributions of each component within CHESS.
The experimental results demonstrate that CHESS outperforms many existing state-of-the-art methods across diverse settings. Specifically, CHESS achieved 71.10% accuracy on the BIRD test set, within 2% of the leading proprietary method Pourreza et al. (2024), while requiring approximately 83% fewer LLM calls. Among the methods using open-source models, CHESS sets a new benchmark with 61.5% accuracy on the BIRD development set. Moreover, we tested CHESS on very large industrial-scale database schema with over 4000 columns. Leveraging the Schema Selector agent, CHESS effectively prunes irrelevant columns, resulting in a 2% increase in accuracy and ×$\times 5$ reduction in token usage. The ablation studies confirm the critical role of components like the Information Retriever and Unit Tester in enhancing overall performance, with each contributing substantially to the accuracy and efficiency of the system. Finally, it is shown that CHESS maintains high accuracy comparable to existing methods even under computational constraints, making it suitable for resource-limited deployments while addressing data privacy concerns.
Contributions. Concretely, here our contributions:
  • Novel Multi-Agent Framework: We propose CHESS, a multi-agent framework that significantly improves text-to-SQL performance and efficiency on complex, real-world databases, utilizing four specialized agents: Information Retriever (IR), Schema Selector (SS), Candidate Generator (CG), and Unit Tester (UT).
  • Innovative Unit Test Generation: We introduce a novel natural language unit test generation method for text-to-SQL, enabling the effective evaluation of generated SQL queries during inference.
  • Computationally Efficient Scalable Retrieval Approach: We develop a highly efficient hierarchical retrieval method for extracting relevant entities and context from vast datasets, significantly enhancing SQL prediction accuracy.
  • Schema Selection Protocol applicable to Industrial-Scale Database Schemas: Our schema selection protocol effectively prunes very large industrial-scale schemas, including those with over 4,000 columns. To our knowledge, this is the first method capable of handling such large schemas at this level of complexity.
  • State-of-the-Art Privacy-Preserving Performance: Among the methods using open-source models, we achieve state-of-the-art performance, resulting in a high-performing, privacy-preserving system suitable for industrial deployment.
The remainder of this paper is organized as follows: Section 2 reviews related work, Section 3 details the CHESS framework, Section 4 presents experimental results, and Section 5 concludes with discussions on future work. All codes to reproduce the results reported in this paper are available on our GitHub repository.***https://github.com/ShayanTalaei/CHESS
@@ 2 Related Work
Generating accurate SQL queries from natural language questions, known as text-to-SQL, is an active research area in NLP and database fields. Initial progress involved custom templates (Zelle and Mooney, 1996), requiring significant manual effort. More recent approaches leverage transformer-based, sequence-to-sequence models (Vaswani et al., 2017; Sutskever et al., 2014), which are well-suited for tasks involving sequence generation, including text-to-SQL (Qin et al., 2022). Initial sequence-to-sequence models, such as IRNet (Guo et al., 2019), used bidirectional LSTM architectures with self-attention for encoding queries and database schemas. To better integrate schema information, models like RAT-SQL (Wang et al., 2019) and RASAT (Qi et al., 2022) incorporated relation-aware self-attention, while SADGA (Cai et al., 2021) and LGESQL (Cao et al., 2021) employed graph neural networks for schema-query relationships. Despite these advancements, sequence-to-sequence models still fall short of human-level performance, with none achieving over 80% execution accuracy on the Spider hold-out test set (Yu et al., 2018).
@@
Alongside the widespread adoption of LLMs across various NLP domains, the text-to-SQL field has similarly benefited from recent methodological innovations with LLMs to enhance performance. Early approaches (Rajkumar et al., 2022), leveraged the zero-shot in-context learning capabilities of LLMs for SQL generation. Building on this, subsequent models including DIN-SQL (Pourreza and Rafiei, 2024a), DAIL-SQL (Gao et al., 2023), MAC-SQL (Wang et al., 2023), and C3 (Dong et al., 2023) have enhanced LLM performance through task decomposition. In addition to in-context learning, proposals in DAIL-SQL (Gao et al., 2023), DTS-SQL (Pourreza and Rafiei, 2024b), and CodeS (Li et al., 2024a) have sought to elevate the capabilities of open-source LLMs through supervised fine-tuning, aiming to rival or exceed their larger, proprietary counterparts. However, the most notable performance gains have been observed in proprietary LLMs utilizing in-context learning methods (Li et al., 2024b). Unlike previous efforts, this paper introduces a hybrid approach that combines both in-context learning and supervised fine-tuning to further enhance performance. Moreover, we propose novel methods to integrate contextual data such as database values and database catalog into the text-to-SQL pipeline, leveraging a rich yet often overlooked source of information. Moreover we also propose natural language generation for SQL query generation to enhance the reliability of the text-to-SQL systems. In contrast to most previous works, the Distillery method (Maamari et al., 2024) demonstrates that the latest versions of LLMs can effectively handle database schema information with up to 200 columns within their prompt, eliminating the need for a separate schema linking step that could introduce errors into the pipeline. In our study, we confirm that for benchmarks like BIRD (Li et al., 2024b), where the schema contains approximately 100 columns, schema linking becomes unnecessary. However, for larger schemas with around 4000 columns, performance degradation still occurs due to the increased prompt size, making a separate schema linking step necessary.
Independently, but concurrently with our work, Large Language Monkeys (Brown et al., 2024), Archon Saad-Falcon et al. (2024), MCS-SQL (Lee et al., 2024), and CHASE-SQL (Pourreza et al., 2024) introduced methods that rely on generating a large number of candidate responses for given questions at inference time. The latter two methods, designed specifically for the text-to-SQL domain, use a selection algorithm to compare and select the best candidate responses. However, unlike these approaches that require numerous LLM calls, we propose a framework of multiple agents that can be used adaptively based on the available computational budget. This framework can either generate multiple candidates and filter them using unit tests, or, in cases where minimizing LLM calls is essential, it allows the combination of different agents to achieve accurate results while staying within budget constraints.
Figure 2: CHESS multi-agent framework for Text-to-SQL task including four agents, Information Retriever (IR), Schema Selector (SS), Candidate Generation (CG), and Unit Tester (UT).
@@ 3 Methodology
“CHESS: Contextual Harnessing for Efficient SQL Synthesis” is a multi-agent framework consisting of four agents, each equipped with specialized tools to achieve specific objectives. Below, we provide a detailed explanation of the functionality and tools of each agent. See Figure 2 for an illustration of overall structure of CHESS.
Real-world deployment of text-to-SQL systems must consider various constraints, such as the reasoning capabilities and sampling budgets of large language models (LLMs), as well as data privacy concerns. CHESS is designed as a configurable framework that can adapt to different deployment scenarios while maintaining high performance. In subsection 4.2, we discuss various deployment settings and how CHESS can be customized to meet performance requirements while adhering to these constraints.
3.1 Agents: Descriptions and Tools
In this subsection, we introduce the core agents of the CHESS framework and their respective tools, each specifically designed to tackle the unique challenges of the text-to-SQL task. The four agents, Information Retriever (IR), Schema Selector (SS), Candidate Generator (CG), and Unit Tester (UT), play essential roles in the framework, leveraging specialized tools to perform their assigned functions. Below, we provide a detailed explanation of each agent’s role and the tools they employ to enable efficient and high-performance SQL query generation. Implementation details and prompt templates are provided in Appendix A and C.
Information Retriever (IR) The information retriever agent equipped with the tools, extract_keywords, entity_retriever, and context_retriever, gathers relevant information related to the input, including entities mentioned in the question and the contextual information provided in the database catalog. The tools are described as follows:
extract_keywords. To search for the similar values in the database and schema description, the agent needs to extract the main keywords from the natural language question. This tool uses a few-shot LLM call to extract the primary keywords and key phrases from the input.
@@ 3.1 Agents: Descriptions and Tools
retrieve_entity. From the list of keywords extracted from the question, some may correspond to entities present in the database. This tool allows the agent to search for similar values in the database and return the most relevant matches, along with their corresponding columns, for each keyword. To assess syntactic similarity between the keywords and the database values, we employ the edit distance similarity metric. Additionally, to enhance retrieval efficiency, we propose a hierarchical retrieval strategy based on Locality Sensitive Hashing (LSH) and semantic (embedding) similarity measures, which we explain in detail in Appendix A. This approach enables the agent to retrieve values that exhibit both syntactic and semantic similarity to the keywords effectively.
@@
retrieve_context. In addition to retrieving values, the IR agent can access the database catalog, which often includes schema metadata, such as column descriptions, extended column names (to resolve abbreviations), and value descriptions. As shown in Figure 1, this information can be useful, and not providing it to the model can lead to suboptimal performance. The IR agent uses this tool to retrieve the most relevant descriptions from the database catalog by querying a vector database of descriptions, constructed during the preprocessing. Retrieval is based on semantic (embedding) similarity, ensuring that the most relevant context is provided to the model.
Schema Selector (SS) The goal of the Schema Selector (SS) agent is to reduce the schema size by selecting only the necessary tables and columns required for generating the SQL query. To achieve this, the SS agent is equipped with three tools, filter_column, select_tables, and select_columns.
filter_column Databases often contain hundreds of columns, many of which may be semantically irrelevant to the user’s query. Inspired by this observation, we design this tool which takes a column name and the question as input and determines whether the column is relevant. While this task could be perform using embedding-based similarity methods, we opt for a relatively inexpensive LLM to ensure high accuracy.
Although this tool is effective at filtering out irrelevant columns, identifying relevant schema elements accurately often requires processing multiple schema items together. The following tools complement this process by handling such scenarios.
select_tables This tool takes a sub-schema of the database and the question as input, and through LLM prompting, returns the tables that are necessary for answering the query.
select_columns To further narrow down the relevant schema items, the select_columns tool can be used. By inputting a sub-schema and the question, the agent retrieves only the necessary columns.
The SS agent can utilize these tools either individually or in combination, depending on its configuration. However, its actions involve a trade-off between precision and recall when selecting relevant schema items. Notably, when dealing with extremely large databases—containing over 4,000 columns—this agent can significantly reduce irrelevant information. A detailed discussion of the importance of SS agent is presented in Section 4.2.2.
Candidate Generator (CG) The Candidate Generator (CG) is responsible for synthesizing SQL query that answers the question asked from the database. To accomplish this, the CG agent first calls the generate_candidate_query tool to generate candidate queries. It then executes these candidates on the database, checks the result, and identifies any faulty queries (those containing syntactic errors or empty result). For each faulty candidate, the agent repeatedly attempts to revise the query, until the issue is resolved or a maximum number of allowed revisions is reached.
generate_candidate_query This tool generates a single candidate query that answers the question. It takes the question, the schema, and the context (entities and descriptions) and prompts an LLM to follow a multi-step reasoning guideline to write a candidate SQL query.
revise Sometimes, the generated candidate queries may contain syntax errors or produce empty result. In such cases, the agent identifies the issue by executing the queries on the database and then uses this tool to fix them. This tool takes the question, schema, context, the faulty query and a description of the issue as input. It then prompts an LLM with this information, asking it revise the query. Note that the issue description (or execution log) is critical in guiding the model, as it provides a direct signal of the query’s failure point.
Unit Tester (UT) The Unit Tester (UT) agent is responsible for selecting the most accurate SQL query from the pool of candidates generated by the CG agent. UT identifies the best candidate by 1) generating multiple unit tests that highlight differences between the candidate queries and 2) evaluating the candidates against these unit tests. It then assigns a score to each query based on the number of unit tests it passes, selecting the top-scoring candidate as the final SQL query for the given question.
generate_unit_test This tool prompts an LLM to generate $k$ unit tests, where $k$ is an input parameter, designed such that only the correct SQL query can pass each of them. The prompt is carefully constructed to produce high-quality unit tests that highlight the semantic differences between the candidate queries. Detailed prompt used to generate unit tests is provided in Appendix C.
evaluate After generating the unit tests, the UT agent evaluates the candidate queries against them. This tool takes multiple candidate queries and a single unit test as input, prompting an LLM to reason through each candidate and determine whether it passes the unit test. While this tool could be implemented to evaluate a single candidate against multiple unit tests simultaneously, our experiments in Section 4.2.1 have shown that the current approach, evaluating multiple candidates against one unit test at a time, yields better results.